### v1

# Analysis of Total Variation Minimization for Clustered Federated Learning

Abstract
A key challenge in federated learning applications is the statistical heterogeneity of local datasets. Clustered federated learning addresses this challenge by identifying clusters of local datasets that are approximately homogeneous. One recent approach to clustered federated learning is generalized total variation minimization (GTVMin). This approach requires a similarity graph which can be obtained by domain expertise or in a data-driven fashion via graph learning techniques. Under a widely applicable clustering assumption, we derive an upper bound the deviation between GTVMin solutions and their cluster-wise averages. This bound provides valuable insights into the effectiveness and robustness of GTVMin in addressing statistical heterogeneity within federated learning environments.
Index Terms:
machine learning, federated learning, distributed algorithms, convex optimization, complex networks
I Introduction
Federated Learning (FL) is an umbrella term for distributed optimization techniques to train machine learning (ML) models from decentralized collections of local datasets [6, 7, 8, 9, 10]. The most basic variant of FL trains a single global model in a distributed fashion from local datasets. However, some FL applications require to train separate (personalized) models for each local dataset [11, 12, 13].
To train high-dimensional personalized models from (relatively) small local datasets, we can exploit the information provided by a similarity graph. The nodes of the similarity graph carry local datasets and corresponding local models. Undirected weighted edges in the similarity graph represent pairwise similarities between the statistical properties of local datasets. One natural approach to exploit the information provided by a similarity graph is generalized total variation minimization (GTVMin) [14]. GTVMin couples the training of personalized models via penalizing the variation of the model parameters across the edges of the similarity graph.
We obtain different instances of GTVMin by using different measures of the variation of model parameters across the edges of the similarity graph. Two well-known special cases of GTVMin are “MOCHA” [10] and network Lasso [15]. Our own recent work studies a GTVMin variant that can handle networks of different (including non-parametric) personalized models [16].
GTVMin is computationally attractive as it can be solved with scalable distributed optimization methods such as stochastic gradient descent or primal-dual methods [17, 18]. Moreover, using a suitable choice for the similarity graph, GTVMin is able to capture the intrinsic cluster structure of local datasets [14].
Contribution. We analyze the cluster structure of GTVMin instances that use the squared Euclidean norm to measure the variation of personalized model parameters. In particular, we provide an upper bound on the cluster-wise variability of model parameters learnt by GTVMin. This analysis complements our own recent work on the cluster structure for GTVMin instances that use variation measures being a norm [14].
Outline. Section II formulates the problem of clustered FL (CFL) for distributed collections of data via generalized total variation minimization (GTVMin) over a similarity graph. Section III contains our main result which is an upper bound on the variation of learnt model parameters across nodes in the same cluster.
II Problem Formulation
In what follows, we develop a precise mathematical formulation of clustered FL (CFL) over networks. Section II-A formulates the problem of learning personalized models for data generators that form clusters. Section II-B defines the concept of a similarity graph that provides information about the pairwise similarities between data generators. Section II-C then uses the similarity graph to formulate GTVMin. Our main result is an upper bound on the cluster-wise variability of local model parameters delivered by GTVMin (see Section III).
II-A Clustered Federated Learning
We consider a collection of $n$ data generators (or “users”) that we index by $i=1,\ldots,n$. Each data generator $i$ delivers a local (or personal) dataset $\mathcal{D}^{(i)}$. The goal is to train a personalized model $\mathcal{H}^{(i)}$, with model parameters $\mathbf{w}^{(i)}$, for each $i$. The usefulness of a specific choice $\mathbf{w}^{(i)}$ for the model parameters is measured by a non-negative local loss function $L_{i}\left(\mathbf{w}^{(i)}\right)\geq 0$.
The common idea of CFL methods is to pools (or cluster) local datasets with similar statistical properties. We can then train a personalized model using the pooled local datasets of the corresponding cluster. CFL is successful if the data generators actually form clusters, within which they are (approximately) homogeneous statistically. We make this requirement precise in the following clustering assumption.
Assumption 1.
Each data generator belongs to some cluster $\mathcal{C}\subseteq\{1,\ldots,n\}$. There is a cluster-specific choice $\overline{{\bf w}}^{(\mathcal{C})}$ for the local parameters for all $i\in\mathcal{C}$ such that
$\sum_{i\in\mathcal{C}}L_{i}\left(\overline{{\bf w}}^{(\mathcal{C})}\right)\leq\varepsilon^{(\mathcal{C})}.$ (1)
Note that the Assumption 1 might be valid for different choices of clusters. Using larger clusters in Assumption 1 requires a larger value $\varepsilon^{(\mathcal{C})}$ for (1) to hold. Unless stated otherwise, we assume that the $\mathcal{C}$ consists of the nodes $i=1,\ldots,|\mathcal{C}|$.
Example. It is instructive to consider Assumption 1 for the special case of local linear regression. Here, generator $i$ delivers $m_{i}$ data points
$\left({\bf x}^{(i,1)},y^{(i,1)}\right),\ldots,\left({\bf x}^{(m_{i})},y^{(m_{i})}\right),$
which we represent by the label vector ${\bf y}^{(i)}=\big{(}y^{(i,1)},\ldots,y^{(i,m_{i})}\big{)}^{T}$ and feature matrix ${\bf X}^{(i)}:=\big{(}{\bf x}^{(i,1)},\ldots,{\bf x}^{(i,m_{i})}\big{)}^{T}$. We assess local model parameters $\mathbf{w}^{(i)}$ via the local loss function $L_{i}\left(\mathbf{w}^{(i)}\right)=(1/m_{i})\left\|{{\bf y}^{(i)}-{\bf X}^{(i)}\mathbf{w}^{(i)}}\right\|_{2}^{2}$. A sufficient condition for Assumption 1 to hold is
${\bf y}^{(i)}\!=\!{\bf X}^{(i)}\overline{{\bf w}}^{(\mathcal{C})}\!+\!{\bm{\varepsilon}}^{(i)},\mbox{ for all }i\in\mathcal{C}.$ (2)
The constant $\varepsilon^{(\mathcal{C})}$ in (1) is then given by
$\varepsilon^{(\mathcal{C})}=\sum_{i\in\mathcal{C}}(1/m_{i})\left\|{\mathbf{w}^{(i)}}\right\|_{2}^{2}.$ (3)
II-B Similarity Graph
In general, we do not know to which cluster a given data generator $i$ belongs to (see Assumption 1). However, we might still have some information about pair-wise similarities $A_{i,i^{\prime}}$ between any two data generators $i,i^{\prime}$. We represent the pair-wise similarities between data generators by an undirected weighted “similarity graph” $\mathcal{G}=\left(\mathcal{V},\mathcal{E}\right)$.
The nodes $\mathcal{V}=\{1,\ldots,n\}$ of this similarity graph $\mathcal{G}$ are the data generators $i=1,\ldots,n$. An undirected edge $\{i,i^{\prime}\}\!\in\!\mathcal{E}$ between two different nodes (data generators) $i,i^{\prime}\in\mathcal{V}$ indicates that they generate data with similar statistical properties. We quantify the extend of this similarity by a positive edge weight $A_{i,i^{\prime}}\!>\!0$. Figure 1 depicts an example of a similarity graph that consists of three clusters.
[…]
Figure 1: Example of a similarity graph whose nodes $i\in\mathcal{V}$ represent data generators and corresponding personalized models. Each personalized model is parametrized by local model parameters $\mathbf{w}^{(i)}$. Two nodes $i,i^{\prime}\in\mathcal{V}$ are connected by an edge $\{i,i^{\prime}\}\in\mathcal{E}$ if the corresponding data generators are statistically similar. The extend of similarity is quantified by a positive edge weight $A_{i,i^{\prime}}$ (indicated by the thickness).
Ultimately, the similarity graph is a design choice for FL methods. This design choice might be guided by domain expertise: data generators being weather stations might be statistically similar if they are located nearby [19]. Instead of domain expertise, we can also use established statistical tests to determine if two local datasets are obtained from a similar (identical) distribution [20].
We can also obtain similarity measures for data generators via estimators for the divergence between probability distributions [21]. The edge weight $A_{i,i^{\prime}}$ can also be determined by a two-step procedure: (i) map each local dataset $\mathcal{D}^{(i)}$ to a vector representation ${\bf z}^{(i)}$ and (ii) evaluate the Euclidean distance between the representations ${\bf z}^{(i)}$ and ${\bf z}^{(i^{\prime})}$.
Ideally, the connectivity of a similarity graph reflects the cluster structure of data generators: Nodes $i\in\mathcal{C}$ in the same cluster (see Assumption 1) should be connected via many edges with large weight. On the other hand, there should only be few boundary (low-weight) edges that connect nodes in- and outside the cluster (see Figure 2).
We measure the internal connectivity of a cluster via the second smallest eigenvalue $\lambda_{2}(\mathcal{C})$ of the Laplacian matrix ${\bf L}^{(\mathcal{C})}$, which is associated with the induced sub-graph $\mathcal{G}^{(\mathcal{C})}$. [^1] The induced sub-graph consists of the cluster nodes $\mathcal{C}$ and all edges $\{i,i^{\prime}\}\in\mathcal{E}$ of the similarity graph $\mathcal{G}$ with $i,i^{\prime}\in\mathcal{C}$.
The larger $\lambda_{2}\big{(}\mathcal{C}\big{)}$, the better the connectivity among the nodes in $\mathcal{C}$. While $\lambda_{2}\big{(}\mathcal{C}\big{)}$ describes the intrinsic connectivity of a cluster $\mathcal{C}$, we also need to characterize its connectivity with the other nodes in the similarity graph. To this end, we will use the cluster boundary
$\left|\partial[\right|\mathcal{C}]\!:=\!\sum_{\{i,i^{\prime}\}\in\partial\mathcal{C}}A_{i,i^{\prime}}\mbox{, with }\partial\mathcal{C}\!:=\!\big{\{}\{i,i^{\prime}\}\!\in\!\mathcal{E}:i\in\mathcal{C},i^{\prime}\!\notin\!\mathcal{C}\big{\}}.$ (4)
For a single-node cluster $\mathcal{C}=\{i\}$, the cluster boundary coincides with the node degree, $\left|\partial[\right|\mathcal{C}]=\sum_{i^{\prime}\neq i}A_{i,i^{\prime}}$.
$\mathcal{C}$$\mathbf{w}^{(1)}$$\mathbf{w}^{(2)}$$\mathbf{w}^{(3)}$$\mathbf{w}^{(4)}$$\mathbf{w}^{(5)}$$\partial\mathcal{C}$
Figure 2: The similarity graph for a collection of data generators that include a cluster $\mathcal{C}$. Ideally, a similarity graph contains many edges between nodes in $\mathcal{C}$ but only few boundary edges (see (4)) between nodes in- and outside $\mathcal{C}$.
II-C Generalized Total Variation Minimization
The goal of CFL is to train a local (or personalized) model $\mathcal{H}^{(i)}$ for each data generator (or user) $i$. Our focus is on local models that are parametrized by vectors $\mathbf{w}^{(i)}\in\mathbb{R}^{d}$, for $i=1,\ldots,n$. The usefulness of a specific choice for the parameters $\mathbf{w}^{(i)}$ is measured by a local loss function $L_{i}\left(\mathbf{w}^{(i)}\right)$, for $i=1,\ldots,n$.
In principle, we could learn $\mathbf{w}^{(i)}$ by minimizing $L_{i}\left(\mathbf{w}^{(i)}\right)$, i.e., implementing a separate empirical risk minimization for each node $i\in\mathcal{V}$. However, this approach fails for a high-dimensional local model $\mathcal{H}^{(i)}$ as they typically require much more training data than provided by the local dataset $\mathcal{D}^{(i)}$.
We can use the similarity graph to regularize the training of personalized models. In particular, we penalize local model parameters that result in a large total variation (TV):
$\sum_{\{i,i^{\prime}\}\in\mathcal{E}}A_{i,i^{\prime}}\left\|{{\bf w}^{(i)}-{\bf w}^{(i^{\prime})}}\right\|_{2}^{2}$ $={\bf w}^{T}\big{(}{\bf L}^{(\mathcal{G})}\otimes\mathbf{I}\big{)}{\bf w}$
[…]
GTVMin optimally balances between the average local loss (training errors) incurred by local model parameters and their TV (II-C),
$\big{\{}\widehat{{\bf w}}^{(i)}\big{\}}_{i=1}^{n}$ $\in\operatorname*{argmin}_{{\bf w}^{(i)}}\bigg{[}\sum_{i\in\mathcal{V}}L_{i}\left({\bf w}^{(i)}\right)+$
$\hskip 28.45274pt\alpha\sum_{\{i,i^{\prime}\}\in\mathcal{E}}A_{i,i^{\prime}}\left\|{{\bf w}^{(i)}-{\bf w}^{(i^{\prime})}}\right\|_{2}^{2}\bigg{]}.$ (6)
If the similarity graph reflects the cluster structure of data generators (see Figure 2), GTVMin (II-C) enforces the learnt parameter vectors $\big{\{}\widehat{{\bf w}}^{(i)}\big{\}}_{i=1}^{n}$ to be approximately constant at cluster nodes (see Assumption 1). Note, however, that GTVMin (II-C) does not require the knowledge about the actual clusters but only the similarity graph.
We can interpret the (weighted edges of the) similarity graph as “hints” offered to GTVMin. If there are enough (and correct) hints, GTVMin recovers the actual cluster structure of data generators, i.e., the learnt model parameters (II-C) are approximately identical for all nodes $i$ in the same cluster $\mathcal{C}$.
Our main result is an upper bound on deviation
$\widetilde{{\bf w}}^{(i)}:=\widehat{\mathbf{w}}^{(i)}-(1/|\mathcal{C}|)\sum_{i^{\prime}\in\mathcal{C}}\widehat{\mathbf{w}}^{(i^{\prime})}$ (7)
between the learnt parameters $\widehat{{\bf w}}^{(i)}$ at cluster nodes $i\in\mathcal{C}$ and their cluster-wise average. This upper bound will involve two key characteristics of a cluster $\mathcal{C}\subseteq\mathcal{V}$: the boundary (4) and the second-smallest eigenvalue $\lambda_{2}(\mathcal{C})$ of the graph Laplacian ${\bf L}^{(\mathcal{C})}$. This eigenvalue allows to lower bound the variation of local model parameters across $\mathcal{C}$,
[…]
$\lambda_{2}(\mathcal{C})\sum_{i\in\mathcal{C}}\left\|{{\bf w}^{(i)}\!-\!{\rm avg}^{(\mathcal{C})}\{\mathbf{w}^{(i)}\}}\right\|_{2}^{2}.$ (11)
Here, ${\rm avg}^{(\mathcal{C})}\{\mathbf{w}^{(i)}\}:=(1/|\mathcal{C}|)\sum_{i\in\mathcal{C}}{\bf w}^{(i)}$ is the average of the local model parameters of cluster nodes $i\in\mathcal{C}$. The bound (10) can be verified via the Courant–Fischer–Weyl min-max characterization [22, Thm. 8.1.2.] for the eigenvalues of the psd matrix ${\bf L}^{(\mathcal{C})}\otimes\mathbf{I}$.
The RHS in (10) has a particular geometric interpretation: It is the squared Euclidean norm of the projection $\mathbf{P}_{\mathcal{S}^{\perp}}{\bf w}^{(\mathcal{C})}$ of the stacked model parameters ${\rm stack}^{(\mathcal{C})}\big{\{}\mathbf{w}^{(i)}\big{\}}\in\mathbb{R}^{d\cdot|\mathcal{C}|}$ on the orthogonal complement $\mathcal{S}^{\perp}$ of the subspace
$\mathcal{S}:=\bigg{\{}\big{(}{\bf c}^{T},\ldots,{\bf c}^{T}\big{)}^{T}\mbox{ for some }{\bf c}\in\mathbb{R}^{d}\bigg{\}}\subseteq\mathbb{R}^{d\cdot|\mathcal{C}|}.$ (12)
The subspace (12) can also be used to decompose the estimation error $\Delta{\bf w}^{(i)}:=\widehat{\mathbf{w}}^{(i)}-\overline{{\bf w}}^{(\mathcal{C})}$ of GTVMin (II-C). Indeed, by stacking the estimation error into a vector $\Delta{\bf w}={\rm stack}^{(\mathcal{C})}\big{\{}\Delta{\bf w}^{(i)}\big{\}}$, we have the orthogonal decomposition
$\Delta{\bf w}=\mathbf{P}_{\mathcal{S}}\Delta{\bf w}+\mathbf{P}_{\mathcal{S}^{\perp}}\Delta{\bf w}.$ (13)
We can evaluate the components in (13) as
$\mathbf{P}_{\mathcal{S}}\Delta{\bf w}={\rm stack}^{(\mathcal{C})}\big{\{}{\rm avg}^{(\mathcal{C})}\{\widehat{\mathbf{w}}^{(i)}\}-\overline{{\bf w}}^{(\mathcal{C})}\big{\}},$ (14)
and
$\mathbf{P}_{\mathcal{S}^{\perp}}\Delta{\bf w}={\rm stack}^{(\mathcal{C})}\big{\{}\widehat{\mathbf{w}}^{(i)}-{\rm avg}^{(\mathcal{C})}\{\widehat{\mathbf{w}}^{(i)}\}\big{\}},$ (15)
III Main Result
Intuitively, we expect GTVMin (II-C) to deliver (approximately) identical model parameters $\mathbf{w}^{(i)}$ for any cluster $\mathcal{C}$ that contains many internal edges but only few boundary edges. Using $\lambda_{2}\big{(}{\bf L}^{(\mathcal{C})}\big{)}$ as a measure for internal connectivity of $\mathcal{C}$ and the boundary measure $\left|\partial[\right|\mathcal{C}]$ (see (4)) we can make this intuition precise.
Theorem 1.
Consider a similarity graph $\mathcal{G}$ whose nodes $i\in\mathcal{V}$ represent data generators and correspond model parameters $\mathbf{w}^{(i)}$. We learn model parameters $\widehat{\mathbf{w}}^{(i)}$, for each node $i\in\mathcal{V}$, via solving GTVMin (II-C). If there is a cluster $\mathcal{C}\subseteq\mathcal{V}$ satisfying Assumption 1,
$\sum_{i\in\mathcal{C}}\left\|{\widetilde{{\bf w}}^{(i)}}\right\|_{2}^{2}$ $\!\leq\!\frac{1}{\alpha\lambda_{2}\big{(}{\bf L}^{(\mathcal{C})}\big{)}}\bigg{[}\varepsilon^{(\mathcal{C})}\!+\!\alpha\left|\partial[\right|\mathcal{C}]2\bigg{(}\left\|{\overline{{\bf w}}^{(\mathcal{C})}}\right\|_{2}^{2}\!+\!R^{2}\bigg{)}\bigg{]}.$ (16)
Here, $R$ denotes an upper bound on the Euclidean norm $\left\|{\widehat{\mathbf{w}}^{(i)}}\right\|_{2}$ outside the cluster, i.e., $\max_{i\in\mathcal{V}\setminus\mathcal{C}}\left\|{\widehat{\mathbf{w}}^{(i)}}\right\|_{2}\leq R$.
Proof.
See Section IV. ∎
Note that Theorem 1 applies to any choice for the non-negative local loss functions $L_{i}\left(\cdot\right)$, for $i=1,\ldots,n$. In particular, the bound (16) applies to any instance of GTVMin as long as the clustering Assumption 1 holds.
The usefulness of the upper bound (16) depends on the availability of a tight bound $R$ on the norm of learnt model parameters outside the cluster $\mathcal{C}$. Such an upper bound can be found trivially, if the loss functions $L_{i}\left(\cdot\right)$ in (II-C) include an implicit constraint of the form $\left\|{\mathbf{w}^{(i)}}\right\|_{2}\leq R$.
We hasten to add that the bound (16) only controls the deviation (7) of the learnt model parameters $\widehat{\mathbf{w}}^{(i)}$ from their cluster-wise average. This deviation coincides with the component (15) of the error $\widehat{\mathbf{w}}^{(i)}-\overline{{\bf w}}^{(i)}$. The bound (16) does not tell us anything about the other error component (14).
Theorem 1 covers single-model FL [6, 23] as the extreme case where all nodes belong to a single cluster $\mathcal{C}=\mathcal{V}$. Trivially, the cluster boundary then vanishes and the bound (16) specializes to
$\sum_{i\in\mathcal{C}}\left\|{\widetilde{{\bf w}}^{(i)}}\right\|_{2}^{2}\!\leq\!\frac{\varepsilon^{(\mathcal{C})}}{\alpha\lambda_{2}\big{(}{\bf L}^{(\mathcal{C})}\big{)}}.$
Thus, for the single-model setting (where $\mathcal{C}=\mathcal{V}$) the error component (7) can be made arbitrarily small by choosing the GTVMin parameter $\alpha$ sufficiently large.
IV Proof of Theorem 1
We verify (16) via a proof by contradiction, i.e., we show that if (16) would not hold, then $\widehat{\mathbf{w}}^{(i)}$ cannot be a solution to (II-C). To this end, we decompose the objective function in GTVMin (II-C) as follows:
$f({\bf w})$ $=$
$\underbrace{\sum_{i\in\mathcal{C}}L_{i}\left(\mathbf{w}^{(i)}\right)\!+\!\alpha\sum_{\begin{subarray}{c}\{i,i^{\prime}\}\in\mathcal{E}\\ i\in\mathcal{C}\end{subarray}}A_{i,i^{\prime}}\left\|{{\bf w}^{(i)}\!-\!{\bf w}^{(i^{\prime})}}\right\|_{2}^{2}}_{=:f^{\prime}\big{(}{\bf w}\big{)}}$ (19)
$+f^{\prime\prime}\big{(}{\bf w}\big{)}.$ (20)
Here, we used the stacked local model parameter ${\bf w}={\rm stack}\big{\{}\mathbf{w}^{(i)}\big{\}}_{i=1}^{n}\in\mathbb{R}^{d\cdot n}$. Note that only the first component $f^{\prime}$ in (IV) depends on the local model parameters $\mathbf{w}^{(i)}$ at the cluster nodes $i\in\mathcal{C}$.
Let us introduce the shorthand $f^{\prime}\big{(}{\bf w}^{(i)}\big{)}$ for the function obtained from $f^{\prime}({\bf w})$ for varying $\mathbf{w}^{(i)}$, $i\in\mathcal{C}$, but fixing $\mathbf{w}^{(i)}:=\widehat{\mathbf{w}}^{(i)}$ for $i\notin\mathcal{C}$. We verify the bound (16) by showing that if it does not hold, the local model parameters $\overline{{\bf w}}^{(i)}:=\overline{{\bf w}}^{(\mathcal{C})}$, for $i\in\mathcal{C}$, results in a smaller value $f^{\prime}\big{(}\overline{{\bf w}}^{(i)}\big{)}<f^{\prime}\big{(}\widehat{{\bf w}}^{(i)}\big{)}$ than the choice $\widehat{\mathbf{w}}^{(i)}$, for $i\in\mathcal{C}$. This would contradict the fact that $\widehat{{\bf w}}^{(i)}$ is a solution to (II-C).
Then, note that
$f^{\prime}\big{(}\overline{{\bf w}}^{(i)}\big{)}$ $=\sum_{i\in\mathcal{C}}L_{i}\left(\overline{{\bf w}}^{(i)}\right)$
+$+\sum_{\begin{subarray}{c}\{i,i^{\prime}\}\in\mathcal{E}\\ i,i^{\prime}\in\mathcal{C}\end{subarray}}\alpha A_{i,i^{\prime}}\left\|{\overline{{\bf w}}^{(\mathcal{C})}\!-\!\overline{{\bf w}}^{(\mathcal{C})}}\right\|_{2}^{2}+\sum_{\begin{subarray}{c}\{i,i^{\prime}\}\in\mathcal{E}\\ i\in\mathcal{C},i^{\prime}\notin\mathcal{C}\end{subarray}}\alpha A_{i,i^{\prime}}\left\|{\overline{{\bf w}}^{(\mathcal{C})}\!-\!\widehat{{\bf w}}^{(i^{\prime})}}\right\|_{2}^{2}$ (25)
(1)${{\leq}}\varepsilon^{(\mathcal{C})}+\alpha\sum_{\begin{subarray}{c}\{i,i^{\prime}\}\in\mathcal{E}\\ i\in\mathcal{C},i^{\prime}\notin\mathcal{C}\end{subarray}}A_{i,i^{\prime}}\left\|{\overline{{\bf w}}^{(\mathcal{C})}\!-\!\widehat{{\bf w}}^{(i^{\prime})}}\right\|_{2}^{2}$ (28)
$\stackrel{{\scriptstyle(a)}}{{\leq}}\varepsilon^{(\mathcal{C})}+\alpha\sum_{\begin{subarray}{c}\{i,i^{\prime}\}\in\mathcal{E}\\ i\in\mathcal{C},i^{\prime}\notin\mathcal{C}\end{subarray}}A_{i,i^{\prime}}2\bigg{(}\left\|{\overline{{\bf w}}^{(\mathcal{C})}}\right\|_{2}^{2}+\left\|{\widehat{{\bf w}}^{(i^{\prime})}}\right\|_{2}^{2}\bigg{)}$ (31)
$\leq\varepsilon^{(\mathcal{C})}+\alpha\left|\partial[\right|\mathcal{C}]2\bigg{(}\left\|{\overline{{\bf w}}^{(\mathcal{C})}}\right\|_{2}^{2}\!+\!R^{2}\bigg{)}.$ (32)
Step $(a)$ uses the inequality $\left\|{{\bf u}\!+\!{\bf v}}\right\|_{2}^{2}\leq 2\big{(}\left\|{{\bf u}}\right\|_{2}^{2}\!+\!\left\|{{\bf v}}\right\|_{2}^{2}\big{)}$ which is valid for any two vectors ${\bf u},{\bf v}\in\mathbb{R}^{d}$.
On the other hand,
[…]
(10)${{\geq}}\alpha\lambda_{2}\big{(}{\bf L}^{(\mathcal{C})}\big{)}\sum_{i\in\mathcal{C}}\left\|{\widetilde{{\bf w}}^{(i)}}\right\|_{2}^{2}.$ (33)
If the bound (16) would not hold, then by (IV) and (32) we would obtain $f^{\prime}\big{(}\widehat{{\bf w}}^{(i)}\big{)}>f^{\prime}\big{(}\overline{{\bf w}}^{(i)}\big{)}$, which contradicts the fact that $\widehat{{\bf w}}^{(i)}$ solves (II-C).